# IncEventGS: Pose-Free Gaussian Splatting from a Single Event Camera

### v1

@@ Abstract
Implicit neural representation and explicit 3D Gaussian Splatting (3D-GS) for novel view synthesis have achieved remarkable progress with frame-based camera (e.g. RGB and RGB-D cameras) recently. Compared to frame-based camera, a novel type of bio-inspired visual sensor, i.e. event camera, has demonstrated advantages in high temporal resolution, high dynamic range, low power consumption and low latency. Due to its unique asynchronous and irregular data capturing process, limited work has been proposed to apply neural representation or 3D Gaussian splatting for an event camera. In this work, we present IncEventGS, an incremental 3D Gaussian Splatting reconstruction algorithm with a single event camera. To recover the 3D scene representation incrementally, we exploit the tracking and mapping paradigm of conventional SLAM pipelines for IncEventGS. Given the incoming event stream, the tracker firstly estimates an initial camera motion based on prior reconstructed 3D-GS scene representation. The mapper then jointly refines both the 3D scene representation and camera motion based on the previously estimated motion trajectory from the tracker. The experimental results demonstrate that IncEventGS delivers superior performance compared to prior NeRF-based methods and other related baselines, even we do not have the ground-truth camera poses. Furthermore, our method can also deliver better performance compared to state-of-the-art event visual odometry methods in terms of camera motion estimation. Code is publicly available at: https://github.com/wu-cvgl/IncEventGS.
@@ 1 Introduction
Acquiring accurate 3D scene representations from 2D images has long been a challenging problem in computer vision. Serving as a fundamental component in various applications such as virtual/augmented reality and robotics navigation etc., substantial efforts have been dedicated to addressing this challenge over the last few decades. Among those pioneering works, Neural Radiance Fields (NeRF) Mildenhall et al. (2021) and 3D Gaussian Splatting (3D-GS) Kerbl et al. (2023), stand out for its utilization of differentiable rendering technique, and have garnered significant attention due to its capability to recover high-quality 3D scene representation from 2D images. Commonly used sensors for 3D scene reconstruction are usually frame-based cameras, such as the RGB and RGB-D cameras. They usually capture full-brightness intensity images within a short exposure time at a regular frequency. Due to the characteristic of the data capturing process, they often suffer from motion blur or fail to capture accurate and informative intensity information under extreme brightness or darkness in the environment, which would further affect the performance of downstream applications.
@@
The event camera, a bio-inspired sensor, has gained significant attention in recent years for its potential to address the limitations of frame-based cameras under challenging conditions. Unlike conventional cameras, event cameras record brightness changes asynchronously at each pixel, emitting events when a predefined threshold is surpassed. This unique operation offers several advantages over conventional cameras, in terms of high temporal resolution, high dynamic range, low latency and power consumption. Although event cameras have attractive characteristics for challenging environments, they cannot be directly integrated into existing frame-based 3D reconstruction algorithms that rely on processing dense 2D brightness intensity images, due to its time continuous, sparse, asynchronous and irregular data capturing characteristics.
@@
Several pioneering works have been proposed to exploit event stream Kim et al. (2016); Rebecq et al. (2017); Gallego et al. (2018) to recover the motion trajectory and scene representation. While existing methods deliver impressive performance, they usually exploit 2.5D semi-dense depth maps to represent the 3D scene, and bundle adjustment (BA) is hardly being performed, due to the asynchronous and sparse characteristics of event data stream. Klenk et al. (2024) recently proposes to convert event stream into event voxel grids, and then adapt a previous frame-based deep visual odometry pipeline Teed et al. (2023) for accurate camera motion estimation. As neural radiance fields (NeRF) exhibits impressive scene representation capability recently, Klenk et al. (2022), Hwang et al. (2023), Rudnev et al. (2023a), Low & Lee (2023) and Low & Lee (2023) explore to recover the underlying dense 3D scene NeRF representation from event stream, by assuming ground truth poses are available.
@@
In contrast to those works, we propose IncEventGS, an incremental dense 3D scene reconstruction algorithm from a single event camera, by exploiting Gaussian Splatting as the underlying scene representation. Different from prior event-based NeRF reconstruction methods, IncEventGS does not require any ground truth camera poses, which is more challenging and provides more flexibility for real-world application scenarios. To overcome the challenges brought by unknown poses, IncEventGS adopts the tracking and mapping paradigm of conventional SLAM pipelines Mur-Artal & Tardós (2017). In particular, IncEventGS exploits prior explored and reconstructed 3D scene for camera motion estimation of incoming event stream during the tracking stage. Both the 3D-GS scene representation and camera motions are then jointly optimized (i.e. event-based bundle adjustment) during the mapping stage, for more accurate scene representation and motion estimation. The 3D scene is progressively expanded and densified. Both synthetic and real datasets are used to evaluate our method. The experimental results demonstrate that IncEventGS is able to recover the underlying 3D scene representation and camera motion trajectory accurately. In particular, IncEventGS outperforms prior NeRF-based methods and other related baselines in terms of scene representation recovery, even IncEventGS does not have the ground-truth poses. Furthermore, our method also delivers better camera motion estimation accuracy than a most recent state-of-the-art visual odometry algorithm, in terms of both the Absolute Trajectory Error (ATE) metric. The recovered 3D scene representation can be further used to render novel brightness images. Our main contributions can be summarized as follows:
@@
  • We present an incremental 3D Gaussian Splatting reconstruction algorithm from a single event camera, without requiring the ground truth camera poses;
@@
  • The experimental results on both the synthetic and real datasets demonstrate superior performance of our method over prior NeRF based methods and related baselines in terms of novel view synthesis, and better performance over state-of-the-art event-based visual odometry algorithm in terms of camera motion estimation;
  • Compared to prior methods, we are able to efficiently render high-quality brightness images thanks to the powerful representation capability of 3D Gaussian Splatting.
@@ 2 Related Works
Event-based neural radiance fields. Prior works from Klenk et al. (2022), Hwang et al. (2023) and Rudnev et al. (2023a) propose to exploit event stream to recover the neural radiance fields with known camera motion trajectory. Low & Lee (2023) further improves the reconstruction algorithm to handle the situation with sparse and noisy events under non-uniform motion. The recovered neural radiance fields can then be used to render novel view brightness images. The ground truth poses are usually computed from corresponding brightness images via COLMAP Schonberger & Frahm (2016) or provided by indoor motion capturing system. Recently, Qu et al. (2024) proposed to integrate event measurements into an RGB-D implicit neural SLAM framework and achieve robust performance under the situation with motion blur. Li et al. (2024) also propose to exploit event measurements and a single blurry image to recover the underlying neural 3D scene representation. In contrast to those works, IncEventGS conduct incremental 3D scene reconstruction without requiring any prior ground truth poses, which is more challenging and provides more flexibility for practical application scenarios. The method further exploits 3D Gaussian Splatting as the underlying scene representation, which demonstrates better image rendering quality and efficiency, compared to NeRF-based representation.
@@
3D Gaussian Splatting. With the recent success of NeRF Mildenhall et al. (2021) and its further developments Fridovich-Keil et al. (2022); Müller et al. (2022), novel view synthesis utilizing 3D representations like MLPs, voxel grids, or hash tables has advanced significantly. While these NeRF-inspired models perform admirably, they frequently require long training and rendering times for individual scenes. The introduction of 3D Gaussian Splatting Kerbl et al. (2023) proposes a novel explicit 3D representation to further improve both the training and rendering efficiency. Due to its impressive efficient scene representation capability, several pioneering work have been proposed to exploit 3D-GS for incremental 3D reconstruction. For example, Keetha et al. propose an RGBD-based 3D-GS SLAM Keetha et al. (2024), employing an online tracking and mapping system tailored to the underlying Gaussian representation. Yan et al. implement a coarse-to-fine camera tracking approach based on the sparse selection of Gaussians Yan et al. (2024). Matsuki et al. propose to apply 3D Gaussian Splatting to do incremental 3D reconstruction using a single moving monocular or RGB-D camera Matsuki et al. (2024). Huang et al. exploits ORB-SLAM3 to compute accurate camera poses and feeds it into a 3D-GS algorithm for dense mapping Huang et al. (2024). Fu et al. uses monocular depth estimation with 3D-GS Fu et al. (2023). Yugay et al. combine DROID-SLAM Teed & Deng (2021) based camera tracking with active and inactive 3D-GS sub-maps Yugay et al. (2023). Hu et al. propose a novel depth uncertainty model to ensure the selection of valuable Gaussian primitives during optimization Hu et al. (2024). While those methods deliver impressive performance in terms of 3D scene recovery and motion estimation, they usually assume the usage of frame-based images (i.e. either RGB or RGB-D date). In the contrary, we propose to exploit pure event measurements for incremental 3D-GS reconstruction. Two concurrent work have also tried to exploit 3D Gaussians for event-based reconstruction recently, i.e. EvGGS Wang et al. (2024) and Event3DGS Xiong et al. (2024). However, both of them rely on ground-truth poses for training, and EvGGS further constrained to object reconstruction from 360-degree surrounding views, while ours assumes the camera poses are not available and is more challenging.
@@ 3 Method
The overview of our IncEventGS is shown in Fig. 1. Given only a single event camera, IncEventGS incrementally performs tracking and dense mapping under the framework of 3D Gaussian Splatting, to recover both the camera motion trajectory and 3D scene representation simultaneously. The main insight of IncEventGS is to accumulate incoming event data into chunks and treat each chunk as a special “image”. We associate each chunk with a continuous time trajectory parameterization in the $\mathfrak{se}3$ space. Two close consecutive timestamps (i.e., $t_{k}$ and $t_{k+\Delta t}$, where $\Delta t$ is a small time interval) can be randomly sampled within the chunk and the corresponding event stream can then be integrated into an image ${\mathbf{E}}(x)$. Based on the parameterized trajectory, the corresponding camera poses (i.e., $\mathbf{T}_{k}$, $\mathbf{T}_{k+\Delta t}$) can be computed and the images (i.e., $\hat{\mathbf{I}}_{k}$, $\hat{\mathbf{I}}_{k+\Delta t}$) can be further rendered from the 3D Gaussian Splatting. The synthesized image $\hat{\mathbf{E}}(x)$ can be computed for event loss computation. During tracking, we only optimize for the camera motion trajectory of the newly accumulated event chunk and exploit the recovered trajectory to initialize the dense bundle adjustment (BA) algorithm for the mapping stage. During mapping stage, we continuously densify 3D Gaussians for newly explored areas and prune transparent 3D Gaussians. For computational efficiency consideration, we exploit a sliding window of the latest chunks and perform BA only within this window for both 3D-GS reconstruction and motion trajectory estimation. We will detail each component as follows.
Figure 1: The pipeline of IncEventGS. IncEventGS processes incoming event stream by dividing it into chunks and representing the camera trajectory as a continuous model. It randomly samples two close consecutive timestamps to integrate the corresponding event streams. Two brightness images are rendered from 3D Gaussian distributions at the corresponding poses, and we minimize the log difference between the rendered images and the accumulated event images. During initialization, a pre-trained depth estimation model estimates depth from the rendered images to bootstrap the system.
@@
Following 3D-GS Kerbl et al. (2023), the scene is represented by a set of 3D Gaussian primitives, each of which contains mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$ in the world coordinate, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\mathbf{o}\in\mathbb{R}$, and color $\mathbf{c}\in\mathbb{R}^{3}$. To ensure that the covariance matrix remains positive semi-definite throughout the gradient descent, the covariance $\boldsymbol{\Sigma}$ is parameterized using a scale vector $\mathbf{s}\in\mathbb{R}^{3}$ and rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$:
@@
where $\boldsymbol{\Sigma}^{\prime}\in\mathbb{R}^{3\times 3}$ is the 2D covariance matrix, $\boldsymbol{J}\in\mathbb{R}^{2\times 3}$ is the Jacobian of the affine approximation of the projective transformation. After projecting 3D Gaussians onto the image plane, the color of each pixel is determined by sorting the Gaussians according to their depth and then applying near-to-far ${\alpha}$-blending rendering via the following equation:
@@
The derivations presented above demonstrate that the rendered pixel color, denoted as I in (3), is a function that is differentiable with respect to the learnable attributes of all Gaussians, and the camera poses $\mathbf{T}_{c}$. This facilitates our bundle adjustment formulation, accommodating a set of event chunks and inaccurate camera motion trajectories within the framework of 3D-GS.
@@ 3.2 Event Data Formation Model
An event camera records changes of the brightness as a stream of events asynchronously. Every time a pixel brightness change reaches a contrast threshold (i.e. $|L(\mathbf{x},t_{i}+\delta t)-L(\mathbf{x},t_{i})|\geq C$), the camera will trigger an event $e_{i}=(\mathbf{x},t_{i},p_{i})$, where $p_{i}\in(-1,+1)$ is the polarity of the event, $L(\mathbf{x},t_{i})=\log(\mathbf{I}(\mathbf{x},t_{i}))$ is the brightness logarithm of pixel $\mathbf{x}$ at timestamp $t_{i}$, $C$ is the fixed contrast threshold.
To relate 3D-GS representation with the event stream, we sample two close consecutive timestamps (i.e., $t_{k}$ and $t_{k+\Delta t}$, where $\Delta t$ is a small time interval), and accumulate the real measured events within $\Delta t$ to an image $\mathbf{E}(\mathbf{x})$. The accumulation is defined as:
@@
where $e(\mathbf{x},t_{i},p_{i})$ is the $i^{th}$ event within the defined time interval corresponding to pixel $\mathbf{x}$. The corresponding camera poses $T_{k}$ and $T_{k+\Delta t}$ can be interpolated from the camera motion trajectory parameterization, allowing us to render two grayscale images (i.e. $\hat{\mathbf{I}}_{k}$ and $\hat{\mathbf{I}}_{k+\Delta t}$) from the previously recovered 3D-GS. The synthesized accumulated event image $\hat{\mathbf{E}}$ can then be computed as:
@@
Both in tracking and mapping, inspired by the work of Rudnev et al. (2023b), we segment the current event chunks into $n_{seg}$ equal segments according to the number of events, obtaining $n_{seg}$ timestamps that correspond to the end of each segment. We then randomly select one timestamp from these $n_{seg}$ timestamps to serve as $t_{k+\Delta t}$, and we randomly sample an integer $n_{win}$ between the integer bounds $n_{low}$ and $n_{up}$. The index of $t_{k}$ is equal to the index of $t_{k+\Delta t}$ subtract $n_{win}$. $n_{seg}$, $n_{low}$ and $n_{up}$ are hyperparameters. This sampling strategy enables the model to capture both local and global information.

### v4

@@ Abstract
Implicit neural representation and explicit 3D Gaussian Splatting (3D-GS) for novel view synthesis have achieved remarkable progress with frame-based camera (e.g. RGB and RGB-D cameras) recently. Compared to frame-based camera, a novel type of bio-inspired visual sensor, i.e. event camera, has demonstrated advantages in high temporal resolution, high dynamic range, low power consumption, and low latency, which make it favored for many robotic applications. In this work, we present IncEventGS, an incremental 3D Gaussian Splatting reconstruction algorithm with a single event camera, without the assumption of known camera poses. To recover the 3D scene representation incrementally, we exploit the tracking and mapping paradigm of conventional SLAM pipelines for IncEventGS. Given the incoming event stream, the tracker first estimates an initial camera motion based on prior reconstructed 3D-GS scene representation. The mapper then jointly refines both the 3D scene representation and camera motion based on the previously estimated motion trajectory from the tracker. The experimental results demonstrate that IncEventGS delivers superior performance compared to prior NeRF-based methods and other related baselines, even if we do not have the ground-truth camera poses. Furthermore, our method can also deliver better performance compared to state-of-the-art event visual odometry methods in terms of camera motion estimation.
@@ 1 Introduction
Reconstructing accurate 3D scene representations from 2D images has been a long-standing challenge in computer vision and robotics, driving substantial efforts over the past few decades. Among those pioneering works, Neural Radiance Fields (NeRF) [19] and 3D Gaussian Splatting (3D-GS) [11], stand out for their utilization of differentiable rendering techniques, and have garnered significant attention due to their capability to recover high-quality 3D scene representation from 2D images. Commonly used sensors for 3D scene reconstruction are usually frame-based cameras, such as the RGB and RGB-D cameras. They usually capture full-brightness intensity images within a short exposure time at a regular frequency. Due to the characteristic of this data-capturing process, they often suffer from motion blur or fail to capture accurate and informative intensity information under fast motion and low-light conditions, which would further affect the performance of downstream applications.
@@
The event camera, a bio-inspired sensor, has gained significant attention in recent years for its potential to address the limitations of frame-based cameras under challenging conditions. Unlike conventional cameras, event cameras record brightness changes asynchronously at each pixel, emitting events when a predefined threshold is surpassed. This unique operation offers several advantages over conventional cameras, in terms of high temporal resolution, high dynamic range, low latency, and power consumption. Although event cameras have attractive characteristics for challenging environments, they cannot be directly integrated into existing frame-based 3D reconstruction algorithms that rely on processing dense 2D brightness intensity images.
@@
Several pioneering works have been proposed to exploit event stream [12, 23, 3] to recover the motion trajectory and scene representation. While existing methods deliver impressive performance, they usually exploit 2.5D semi-dense depth maps to represent the 3D scene, and bundle adjustment (BA) is hardly performed, due to the asynchronous and sparse characteristics of event data stream. Klenk et al. [15] recently proposed to convert event stream into event voxel grids, and then adapt a previous frame-based deep visual odometry pipeline [30] for accurate camera motion estimation. As NeRF exhibited impressive scene representation capability recently, several works [14, 8, 25, 17] explore to recover the underlying dense 3D scene NeRF representation from event stream, by assuming ground-truth poses are available.
@@
In contrast to those works, we propose IncEventGS, an incremental dense 3D scene reconstruction algorithm from a single event camera, by exploiting Gaussian Splatting as the underlying scene representation. Different from prior event-based NeRF reconstruction methods, IncEventGS does not require any ground-truth camera poses, which is more challenging and provides more flexibility for real-world robotic application scenarios. To overcome the challenges brought by unknown poses, IncEventGS adopts the tracking and mapping paradigm of conventional SLAM pipelines [20]. In particular, IncEventGS exploits prior explored and reconstructed 3D scenes for camera motion estimation of incoming event stream during the tracking stage. Both the 3D-GS scene representation and camera motions are then jointly optimized (i.e. event-based bundle adjustment) during the mapping stage, for more accurate scene representation and motion estimation. The 3D scene is progressively expanded and densified. The experimental results on both synthetic and real datasets demonstrate that IncEventGS can recover the underlying 3D scene representation and camera motion trajectory accurately. In particular, IncEventGS outperforms prior NeRF-based methods and other related baselines in terms of scene representation recovery, even without ground-truth poses. Furthermore, our method also delivers better camera motion estimation accuracy than the most recent state-of-the-art visual odometry algorithm, in terms of the Absolute Trajectory Error (ATE) metric. The recovered 3D scene representation can be further used to render novel brightness images. Our main contributions can be summarized as follows:
@@
  • We present an incremental 3D Gaussian Splatting reconstruction algorithm from a single event camera, without requiring the ground-truth camera poses.
  • We propose a novel initialization strategy tailored to the event data stream, which is vital to the success of the algorithm.
@@
  • The experimental results on both the synthetic and real datasets demonstrate the superior performance of our method over prior NeRF-based methods and related baselines in terms of novel view synthesis and better performance over state-of-the-art event-based visual odometry algorithm in terms of camera motion estimation.
@@ 2 Related Works
Event-based Neural Radiance Fields. Prior works [14, 8, 25] propose to exploit event stream to recover the neural radiance fields with known camera motion trajectory. Low et al. [17] further improves the reconstruction algorithm to handle sparse and noisy events under non-uniform motion. The recovered neural radiance fields can then be used to render novel view brightness images. The ground-truth poses are usually computed from corresponding brightness images via COLMAP [26] or provided by the indoor motion-capturing system. Recently, Qu et al. [22] proposed to integrate event measurements into an RGB-D implicit neural SLAM framework and achieved robust performance in motion blur scenarios. Li et al. [16] also proposed exploiting event measurements and a single blurry image to recover the underlying neural 3D scene representation. In contrast to those works, IncEventGS conducts incremental 3D scene reconstruction without requiring any prior ground-truth poses, which is more challenging and provides more flexibility for practical robotic application scenarios. The method further exploits 3D Gaussian Splatting as the underlying scene representation, which demonstrates better image rendering quality and efficiency, compared to the NeRF-based representation.
@@
3D Gaussian Splatting. 3D Gaussian Splatting [11] proposes a novel explicit 3D representation to further improve both the training and rendering efficiency compared to Neural Radiance Fields. Due to its impressive efficient scene representation capability, several pioneering works have been proposed to exploit 3D-GS for incremental 3D reconstruction. For example, Keetha et al. [10] propose an RGBD-based 3D-GS SLAM, employing an online tracking and mapping system tailored to the underlying Gaussian representation. Yan et al. [36] implement a coarse-to-fine camera tracking approach based on the sparse selection of Gaussians. Matsuki et al. [18] propose to apply 3D Gaussian Splatting to do incremental 3D reconstruction using a single moving monocular or RGB-D camera. Huang et al. [7] exploit ORB-SLAM3 to compute accurate camera poses and feed it into a 3D-GS algorithm for dense mapping. Fu et al. [2] use monocular depth estimation with 3D-GS. Yugay et al. [37] combine DROID-SLAM [29] based camera tracking with active and inactive 3D-GS sub-maps. Wang et al. [33] integrate bundle-adujustment and 3DGS [38] to estimated camera trajectory within the exposure time. Hu et al. [6] propose a novel depth uncertainty model to ensure the selection of valuable Gaussian primitives during optimization. While those methods deliver impressive performance in terms of 3D scene recovery and motion estimation, they usually assume the usage of frame-based images (i.e. either RGB or RGB-D date). On the contrary, we propose to exploit pure event measurements for incremental 3D-GS reconstruction. Several concurrent studies have recently explored the use of 3D Gaussians for event-based reconstruction, including EvGGS [31], Event3DGS [35], and E2GS [1]. EvGGS and Event3DGS rely solely on event data, whereas E2GS incorporates both event data and blurry images. However, the key difference between these methods and ours is that they all rely on ground-truth poses.
@@ 3 Method
Figure 1: The pipeline of IncEventGS. IncEventGS processes incoming event stream by dividing it into chunks and representing the camera trajectory as a continuous model. It randomly samples two close consecutive timestamps to integrate the corresponding event streams. Two brightness images are rendered from 3D-GS at the corresponding poses, and we minimize the photometric loss between the synthesized and measured brightness change. During initialization, a pre-trained depth estimation model estimates depth from the rendered images to bootstrap the system.
The overview of our IncEventGS is shown in Fig. 1. Given only a single event camera, IncEventGS incrementally performs tracking and dense mapping under the framework of 3D Gaussian Splatting, to recover both the camera motion trajectory and 3D scene representation simultaneously. Since event data are asynchronous, they cannot be directly integrated with the 3D-GS representation. We therefore process the event data stream into chunks according to a fixed time window. We associate each chunk with a continuous time trajectory parameterization in the $\mathfrak{se}(3)$ space. Two close consecutive timestamps (i.e., $t_{k}$ and $t_{k+\Delta t}$, where $\Delta t$ is a small time interval) can be randomly sampled, and the measured brightness change ${\mathbf{E}}(x)$ can be computed from the corresponding event stream. Based on the parameterized trajectory, the corresponding camera poses (i.e., $\mathbf{T}_{k}$, $\mathbf{T}_{k+\Delta t}$) can be determined and the brightness images (i.e., $\hat{\mathbf{I}}_{k}$, $\hat{\mathbf{I}}_{k+\Delta t}$) can be further rendered from the 3D-GS. The synthesized brightness change $\hat{\mathbf{E}}(x)$ can be computed for event loss computation.
During the tracking stage, we optimize only the camera motion trajectory of the newly accumulated event chunk and exploit the recovered trajectory to initialize the dense bundle adjustment (BA) algorithm for the mapping stage. During the mapping stage, we continuously densify 3D Gaussians for newly explored areas and prune transparent 3D Gaussians. For computational efficiency, we exploit a sliding window of the latest chunks and perform BA only within this window for both 3D-GS reconstruction and motion trajectory estimation. We will detail each component as follows.
@@
Following 3D-GS [11], the scene is represented by a set of 3D Gaussian primitives, each of which contains mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$ in the world coordinate, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\mathbf{o}\in\mathbb{R}$, and color $\mathbf{c}\in\mathbb{R}^{3}$. To ensure that the covariance matrix remains positive semi-definite throughout the gradient descent, the covariance $\boldsymbol{\Sigma}$ is parameterized using a scale vector $\mathbf{s}\in\mathbb{R}^{3}$ and rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$:
@@
where $\boldsymbol{\Sigma}^{\prime}\in\mathbb{R}^{2\times 2}$ is the 2D covariance matrix, $\boldsymbol{J}\in\mathbb{R}^{2\times 3}$ is the Jacobian of the affine approximation of the projective transformation. After projecting 3D Gaussians onto the image plane, the color of each pixel is determined by sorting the Gaussians according to their depth and then applying near-to-far ${\alpha}$-blending rendering via the following equation:
@@
The derivations presented above demonstrate that the rendered pixel color, denoted as I in Eq. (3), is a function that is differentiable with respect to the learnable attributes of 3D-GS, and the camera poses $\mathbf{T}_{c}$. This facilitates our bundle adjustment formulation, accommodating a set of event chunks and inaccurate camera motion trajectories within the framework of 3D-GS.
@@ 3.2 Event Data Formation Model
An event camera records changes in brightness as a stream of events asynchronously. To relate 3D-GS representation with the event stream, we sample two close consecutive timestamps (i.e., $t_{k}$ and $t_{k+\Delta t}$, where $\Delta t$ is a small time interval), and the measured brightness change between $\Delta t$ is:
@@
where $e(\mathbf{x},t_{i},p_{i})$ is the $i^{th}$ event within the defined time interval corresponding to pixel $\mathbf{x}$ and $C$ is the fixed contrast threshold. The corresponding camera poses $T_{k}$ and $T_{k+\Delta t}$ can be interpolated from the camera motion trajectory parameterization, and two corresponding brightness images (i.e. $\hat{\mathbf{I}}_{k}$ and $\hat{\mathbf{I}}_{k+\Delta t}$) can be rendered from 3D-GS. The synthesized brightness change $\hat{\mathbf{E}}$ is modelled as:
@@
Both in tracking and mapping, inspired by the work of [25], we segment the current event chunks into $n_{seg}$ equal segments according to the number of events, obtaining $n_{seg}$ timestamps that correspond to the end of each segment. We then randomly select one timestamp from these $n_{seg}$ timestamps to serve as $t_{k+\Delta t}$, and we randomly sample an integer $n_{win}$ between the integer bounds $n_{low}$ and $n_{up}$. The index of $t_{k}$ is equal to the index of $t_{k+\Delta t}$ subtract $n_{win}$. $n_{seg}$, $n_{low}$ and $n_{up}$ are hyperparameters. This sampling strategy enables the model to capture both local and global information.